# IoT Ground Sensing Systems for Early Wildfire Detection: Technologies, Challenges and Opportunities

### v1

@@ III Data Analytic Algorithms for Wildfire Detection
After collecting these data from EDs, the integration of camera images or sensor data is used to identify any fire or non-fire event. To do this, either the MCUs on the EDs or the computing unit on the central server must use data-analytic algorithms to convert the problem into a mathematical expression. The analytical process of identifying unusual events within IoT applications often aligns with the principles of anomaly detection. In the context of wildfire detection, this is specifically focused on distinguishing the rare occurrences of fires, while non-fire situations happen most of the time. Each anomaly detection technique has its own benefits and drawbacks, which make them more or less suitable for use in IoT projects.
In this section, a summary of early wildfire detection algorithms is provided. These algorithms can be classified into two main categories based on the way they obtain their detection criteria: feature detection and pattern detection in time series. The discussion of the use of ML and DL to improve the performance of features or pattern detection for both vision-based and non-vision-based detection is also provided, as the use of ML/DL in IoT projects is gaining more research interests. Although ML/DL are heavily dependent on computers to recognise irregularities through iterative learning of features/patterns, it is essential to remember that these techniques are still based on statistical and mathematical models and can be classified as a part of the features/patterns detection. To make it easier for IoT designers in finding relevant information, we have divided this into an individual item. The following subsections will discuss fire/smoke vision-based detection algorithms for recognising anomalies in features/patterns and how to enhance them using ML/DL, followed by detection algorithms for non-vision-based detection for recognising anomalies in features/patterns and how to improve them using ML/DL.
@@ III-A2 Time Series Pattern Detection
Temporal Wavelet Analysis: In addition to feature detection, more sophisticated image processing techniques can be used to detect patterns in time-series data. Temporal wavelet analysis combines colour and motion analysis to detect fires. Moving objects of fire can sometimes trigger a false alarm, but analysing the flickering property of a fire can help distinguish between real fires and these objects [119, 127, 115, 128, 129, 130]. The flickering frequency of the fire is around 10 Hz [131], and an oscillating frequency of a pixel intensity value greater than 0.5 Hz indicates the presence of the fire [132]. [133] used the Discrete Wavelet Transform (DWT) to extract information on temporal variation from a pixel by placing its intensity history in a two-stage filter bank. And by comparing the simulated results for fire pixels and non-fire pixels, fire pixels can be distinguished by counting the number of zero crossings within a certain number of frames [134]. Although temporal wavelet analysis combines motion and colour analysis to enhance fire detection accuracy, this comes at the cost of increased computational complexity and the need for more temporary memory capacity, leading to higher energy consumption and longer detection latency. Furthermore, smoke and flames can flicker in various ways, and to accurately capture the characteristics of early wildfires, various studies have suggested multiple approaches by integrating temporal wavelet analysis with other techniques in fire detection [135, 136].
@@ III-A3 Machine Learning and Deep Learning Algorithms
Although the use of vision-based technologies has shown great potential in the analysis of satellite images, real-time camera systems on UAVs, and alleviating the workload previously carried out by watchtower observers responsible for detecting occurrences of wildfires [164], it remains susceptible to interference from various environmental factors, such as weather conditions (e.g., cloud cover, rainfall and reflection of sunlight), as well as physical obstructions such as trees and buildings. As a result, the response time of imaging detection can be significantly delayed when camera modules are used on IoT EDs. Besides, if the number of cameras mounted on an IoT ED is limited, the camera is usually required to rotate and scan every hour to cover a full 360 degrees of view, which will increase the time it takes to detect something. Furthermore, these camera systems can be expensive to deploy and consume significant amounts of energy. Therefore, early wildfire detection using environmental sensing data in IoT systems has been explored as a potential alternative.
III-B Detection Algorithm based on Environmental Sensing Data
@@ III-B2 Time Series Detection
An alternative approach to analysing time-series data is the seasonality detection method, such as seasonal decomposition methods and seasonal forecasting methods. In seasonality detection, the trend or seasonal component captures the overall direction of the data over a long-term period, providing information on whether the time series is increasing, decreasing, or relatively stable. Seasonal decomposition techniques are used to identify patterns in the data that repeat at fixed intervals by isolating the seasonal component so that any recurring patterns in the data can be retrieved. These fluctuations can be considered anomalies or unexpected behaviour in the data [177]. In contrast, seasonality forecasting techniques, such as Holt-Winters and Fourier analysis, also consider seasonality when predicting. However, to enable anomaly detection, users still need to design their own anomaly-checking mechanism. A straightforward detection algorithm is to calculate the mean square error and feed it into a given threshold and confidence intervals to determine if it is an error. In [178], the author evaluated the effectiveness of various IoT air quality prediction models using seasonal forecasting methods and seasonality ML/DL prediction models. The results of the study showed that it is possible to predict seasonality in EDs using low computing power Cortex-M0 or high-performance Cortex A53. However, seasonal anomaly detection models are not widely used in anomaly detection due to the fact that the full data set may not be available if the seasonal patterns last longer than one year. For example, in [179], researchers have shown that Holt-Winters requires an additional 40KB of data to correctly predict CO and the temperature level on an ESP32. Besides, these methods assume that the time series has repeated seasonal and stationary patterns. If the data includes irregular or non-recurring seasonal patterns or if the seasonal patterns change over time, these methods may not be able to accurately capture the underlying structure.
@@ III-B3 Anomaly Detection Using ML and DL
In addition to the detection methods mentioned, the use of ML or DL for anomaly detection has been increasing due to its ability to capture complex relationships and non-linear patterns in data, as well as its adaptability and scalability in dealing with different input features. These ML/DL approaches make it possible to process large or multi-dimensional datasets quickly and easily for point-wise or time-series detection. Moreover, ML techniques can be applied to various components of a network, from IoT devices to network servers. These advantageous characteristics have led to the growing adoption of ML for anomaly detection. Several basic supervised ML algorithms have been used for anomaly detection, such as SVM, K-nearest neighbours (KNN) and decision tree (DT) [180]. SVM uses hyperplanes to classify data into normal or anomaly groups, while KNN determines anomalies based on the Euclidean distance to their nearest-neighbour group. DT uses reduction of entropy values to identify relevant information for anomaly detection. However, these ML techniques require user-labelled data points, which are usually obtained by manual detection. For instance, when it comes to fire detection, users must label the time stream of sensor readings as either fire or non-fire states. The accuracy of these methods is highly dependent on the accuracy of the labelling process. Therefore, some anomaly detection methods are developed using ML clustering techniques, such as K-means and GMMs, and many IoT projects have implemented these approaches, as shown in [89, 181]. Furthermore, some ML forecasting models, such as Autoregressive Integrated Moving Average (ARIMA) and Seasonal Autoregressive Integrated Moving Average (SARIMA), can also be combined with statistical outlier methods to detect anomalies without supervised learning [182]. If the input dimensions of ML are small, the energy consumption of ML can be comparable to that of statistical methods. However, the greatest advantages of ML are its ability to recognise interconnections between different inputs and to detect anomalies. Furthermore, its complexity and computing power requirements are lower than those of the DL methods, making it more suitable for deployment on resource-constrained EDs. Recent IoT ground sensing system projects have widely used ML anomaly detection algorithms, as summarised in Table IV.
@@
Table. IV presents an summary of the anomaly detection algorithms used in recent forest fire detection projects. In general, DL and ML approaches are widely used in these projects. Most of these initiatives employ a threshold limit, based on the z-score, regression, or LOF, for initial filtering. Upon triggering consecutive alerts, the ED transmits data to the server to perform fire detection or performs specific fire detection algorithms on the ED itself. Note that in this section, we have focused on a few anomaly detection techniques that are suitable for implementation on EDs or the gateway for wildfire detection. The selection of these methods was based on their computing power requirements, complexity, and whether they have been previously implemented.
@@ IV-A Sensors
As mentioned in Sec. II, gas/smoke sensors have different characteristics when it comes to wildfire detection. Some sensors have minimal cross-sensitivity, but their operating temperature and humidity ranges are limited. On the other hand, some sensors are highly accurate, but they take time to precisely measure the concentration of the target gases. Additionally, the cost and availability of these sensors, such as energy consumption, maintenance cost and size, are crucial considerations. It is essential to take into account not only the cost of the sensors, but also the operational costs associated with the equipment and needed for their operation and maintenance. A thorough assessment of these elements is necessary when selecting the right sensors for a particular purpose. Sensors can be delicate and need to be calibrated regularly, which can make them costly to use in a wildfire detection system. They may be exposed to extreme environmental conditions and may need to be replaced often. Besides, the practical implementation of these sensors in the field also presents several challenges, including the establishment of a secure mounting site and the protection of the sensors against environmental hazards. These challenges pose obstacles in deploying sensors in remote regions, which are often the areas most vulnerable to wildfires.
@@ IV-B Communication link and Access Layer Control
Typically, IoT systems employ LPWAN protocols for communication between the end nodes and the gateway. These LPWAN communication protocols are characterised by ultra-narrowband or simple modulation/demodulation techniques, enabling low energy consumption during communication. The use of ultra-narrowband signals reduces the chance of packet collision, while wideband signals with distinctive patterns allow for easy differentiation from other signals, facilitating extended-range communication. Key players in the LPWAN domain include LoRaWAN, Sigfox, 6LoWPAN, NBIoT and LTE-M [192]. Although NBIoT and LTE-M operate within licenced spectrum bands, the LoRaWAN together with Sigfox and 6LoWPAN utilise ISM-shared bands. A performance analysis comparing various wireless communication technologies, as shown in Fig. 6, demonstrates that LPWAN offers the longest range and the lowest energy consumption. However, it should be noted that LPWAN technologies exhibit the lowest data throughput. As a result, IoT system designers must prioritise the significance of data and determine which data to transmit and discard. Although most LPWAN technologies support packet fragmentation, this feature can exaggerate the likelihood of packet collisions and the overall packet loss rate within the system.
@@
In wireless communication, packet collision occurs when multiple packets are simultaneously transmitted to a receiver, making it impossible to differentiate between individual packets. To mitigate this issue, an MAC layer is responsible for regulating channel traffic, thereby ensuring successful packet reception. However, the MAC layer in most LPWAN technologies is typically pre-defined, with some employing random access or simple Listen Before Transmit techniques [193, 194]. In contrast, NB-IoT and LTE-M use the same MAC control as cellular networks due to their integration in the cellular communication system. However, optimisation of packet success rates in IoT systems often requires the implementation of a customised MAC control. In [195], a custom Carrier-sense multiple access with collision avoidance (CSMA/CA) MAC layer was implemented in the LoRa communication system to improve the packet success rate. However, the adoption of an in-house MAC layer control can lead to increased energy consumption for end devices due to requirements such as channel listening and synchronisation. An alternative approach is demonstrated in [196], where an in-house CSMA/CA MAC control for LoRa is introduced, resulting in an improvement in energy consumption. Overall, achieving a robust and energy-efficient communication link in IoT systems presents challenges that require thorough consideration of various factors, including gateway coverage influenced by channel loss and modulation schemes, channel interference resulting from traffic load, packet size and packet priorities, as well as quality of service and packet latency. In order to tackle these obstacles, the implementation of a custom MAC layer control emerges as a viable solution for IoT designers.
@@ IV-D Accuracy and Topology
As mentioned above, to minimise energy consumption in the ED, it is recommended to use energy-efficient detection algorithms. In cases where the decision-making process is based solely on the ED, simpler ML algorithms or statistical threshold detection methods can be employed. However, it is important to note that such an approach can reduce the accuracy of the system, leading to potential issues of excessive sensitivity or the generation of numerous false alarms. To enhance the precision of fire detection systems, a viable solution is to integrate multiple detection aspects, such as combining vision-based detection and detection based on environmental monitoring sensors, to establish a remote sensing system. Nonetheless, this solution may incur additional energy consumption in each individual end device. Consequently, it becomes imperative to investigate the trade-off between accuracy and energy consumption. In addition, it is essential to thoroughly analyse the topology of the sensing system, including deciding the best number of environmental monitoring sensor nodes, the correct number of vision-based detection nodes, and the most suitable spacing and angular arrangement between the sensors. To address these topological issues, extensive measurement and analysis must be performed.
@@
The challenges involved in implementing an IoT sensor system for early wildfire detection in remote areas can be illustrated by the trade-off between energy consumption, detection accuracy and delay. One of the main challenges is to conserve energy within the system. This can be achieved by using energy-friendly algorithms on the end devices and by sending decisions directly to the main server. However, this approach may not be able to guarantee high accuracy. Furthermore, ensuring detection accuracy requires a comprehensive sensing system that uses multiple aspects for detection and incorporates cross-validation techniques. However, this approach results in increased energy consumption and delay. Alternatively, in the pursuit of avoiding packet collision and achieving rapid response, only the detection results are transmitted to the server, which may lead to compromised accuracy. Overall, there is a complex trade-off between energy consumption, detection accuracy and delay in implementing an IoT sensor system for wildfire detection in remote areas. There is no single solution that is optimal for all situations. The best approach will depend on several factors, including the size of the area to be monitored, the availability of resources and the risk of wildfires.
V Potential Framework and Possible Extension

### v2

@@ III Data Analytic Algorithms for Wildfire Detection
Once the data collection from EDs is completed, the integration of camera images or sensor data is used to identify any fire or non-fire event. To do this, either the MCUs on the EDs or the computing unit on the central server must use data-analytic algorithms to convert the problem into a mathematical expression. The analytical process of identifying unusual events within IoT applications often aligns with the principles of anomaly detection. In the context of wildfire detection, this is specifically focused on distinguishing the rare occurrences of fires, while non-fire situations happen most of the time. Each anomaly detection technique has its own benefits and drawbacks, which make them more or less suitable for use in IoT projects.
In this section, a summary of early wildfire detection algorithms is provided. These algorithms can be classified into two main categories based on the way they obtain their detection criteria: feature detection and pattern detection in time series. The discussion of the use of ML and DL to improve the performance of features or pattern detection for both vision-based and non-vision-based detection is also provided, as the use of ML/DL in IoT projects is gaining more research interests. Although ML/DL are heavily dependent on computers to recognise irregularities through iterative learning of features/patterns, it is essential to remember that these techniques are still based on statistical and mathematical models and can be classified as a part of the features/patterns detection. To make it easier for IoT designers in finding relevant information, we have divided this into an individual item. The following subsections discuss vision-based fire and smoke detection algorithms that use various techniques, including ML/DL and temporal/pointwise detection. Subsequently, these subsections also explore nonvision-based detection algorithms to identify anomalies either point-wise or temporally.
@@ III-A2 Time Series Pattern Detection
Temporal Wavelet Analysis: In addition to feature detection, more sophisticated image processing techniques can be used to detect patterns in time-series data. Temporal wavelet analysis combines colour and motion analysis to detect fires. Moving objects of fire can sometimes trigger a false alarm, but analysing the flickering property of a fire can help distinguish between real fires and these objects [123, 132, 119, 133, 134, 135]. The fire flickers at a frequency of approximately 10 Hz [136], and a pixel intensity oscillation frequency above 0.5 Hz suggests the presence of fire [137]. [138] used the Discrete Wavelet Transform (DWT) to extract information on temporal variation from a pixel by placing its intensity history in a two-stage filter bank. And by comparing the simulated results for fire pixels and non-fire pixels, fire pixels can be distinguished by counting the number of zero crossings within a certain number of frames [139]. Although temporal wavelet analysis combines motion and colour analysis to enhance fire detection accuracy, this comes at the cost of increased computational complexity and the need for more temporary memory capacity, leading to higher energy consumption and longer detection latency. Furthermore, smoke and flames can flicker in various ways, and to accurately capture the characteristics of early wildfires, various studies have suggested multiple approaches by integrating temporal wavelet analysis with other techniques in fire detection [140, 141].
@@ III-A3 Machine Learning and Deep Learning Algorithms
Table V presents a summary of the vision-based methods utilised by various IoT wildfires detection systems, detailing the advantages and disadvantages of these methods as discussed in this section. Although the use of vision-based technologies has shown great potential in the analysis of satellite images, real-time camera systems on UAVs, and alleviating the workload previously carried out by watchtower observers responsible for detecting occurrences of wildfires [169], it remains susceptible to interference from various environmental factors, such as weather conditions (e.g., cloud cover, rainfall and reflection of sunlight), as well as physical obstructions such as trees and buildings. As a result, the response time to imaging detection can be significantly delayed when camera modules are used on IoT EDs. Besides, if the number of cameras mounted on an IoT ED is limited, the camera is usually required to rotate and scan every hour to cover a full 360 degrees of view, which will increase the time it takes to detect something. Furthermore, these camera systems can be expensive to deploy and consume significant amounts of energy. Therefore, early wildfire detection using environmental sensing data in IoT systems has been explored as a potential alternative.
III-B Environmental Sensing-Based Detection Algorithm
@@
Algorithm | Reference(s) | Pros | Cons |
Feature Detection | [114, 115, 116, 117, 118, 119, 120, 121, 122] | Easy to deploy | Cannot detect temporal |
[125, 126, 127, 128, 129] | smoke pattern | ||
Time Series Detection | [132, 133, 134, 135, 136, 137, 138, 139, 140, 141] | Temporal pattern | Longer |
(Temporal Detection) | can be extracted | processing time | |
DL Neural Networks | [160, 156, 161, 162, 163] | Highly accurate | Resource requirement |
and well-established | and delay response |
@@ III-B2 Time Series Detection
An alternative approach to analysing time-series data is the seasonality detection method, such as seasonal decomposition methods and seasonal forecasting methods. In seasonality detection, the trend or seasonal component captures the overall direction of the data over a long-term period, providing information on whether the time series is increasing, decreasing, or relatively stable. Seasonal decomposition techniques are used to identify patterns in the data that repeat at fixed intervals by isolating the seasonal component so that any recurring patterns in the data can be retrieved. These fluctuations can be considered anomalies or unexpected behaviour in the data [182]. In contrast, seasonality forecasting techniques, such as Holt-Winters and Fourier analysis, also consider seasonality when predicting. However, to enable anomaly detection, users still need to design their own anomaly-checking mechanism. A straightforward detection algorithm is to calculate the mean square error and feed it into a given threshold and confidence intervals to determine if it is an error. In [183], the author evaluated the effectiveness of various IoT air quality prediction models using seasonal forecasting methods and seasonality ML/DL prediction models. The results of the study showed that it is possible to predict seasonality in EDs using low computing power Cortex-M0 or high-performance Cortex A53. However, seasonal anomaly detection models are not widely used in anomaly detection due to the fact that the full data set may not be available if the seasonal patterns last longer than one year. For example, in [184], researchers have shown that Holt-Winters requires an additional 40KB of data to correctly predict CO2 and the temperature level on an ESP32. Besides, these methods assume that the time series has repeated seasonal and stationary patterns. If the data includes irregular or non-recurring seasonal patterns or if the seasonal patterns change over time, these methods may not be able to accurately capture the underlying structure.
@@ III-B3 Anomaly Detection Using ML and DL
In addition to the detection methods mentioned, the use of ML or DL for anomaly detection has been increasing due to its ability to capture complex relationships and non-linear patterns in data, as well as its adaptability and scalability in dealing with different input features. These ML/DL approaches make it possible to process large or multi-dimensional datasets quickly and easily for point-wise or time-series detection. Moreover, ML techniques can be applied to various components of a network, from IoT devices to network servers. These advantageous characteristics have led to the growing adoption of ML for anomaly detection. Several basic supervised ML algorithms have been used for anomaly detection, such as SVM, K-nearest neighbours (KNN) and decision tree (DT) [185]. SVM uses hyperplanes to classify data into normal or anomaly groups, while KNN determines anomalies based on the Euclidean distance to their nearest-neighbour group. DT uses reduction of entropy values to identify relevant information for anomaly detection. However, these ML techniques require user-labelled data points, which are usually obtained by manual detection. For instance, when it comes to fire detection, users must label the time stream of sensor readings as either fire or non-fire states. The accuracy of these methods is highly dependent on the accuracy of the labelling process. Therefore, some anomaly detection methods are developed using ML clustering techniques, such as K-means and GMMs, and many IoT projects have implemented these approaches, as shown in [93, 186]. Furthermore, some ML forecasting models, such as Autoregressive Integrated Moving Average (ARIMA) and Seasonal Autoregressive Integrated Moving Average (SARIMA), can also be combined with statistical outlier methods to detect anomalies without supervised learning [187]. If the input dimensions of ML are small, the energy consumption of ML can be comparable to that of statistical methods. However, the greatest advantages of ML are its ability to recognise interconnections between different inputs and to detect anomalies. Furthermore, its complexity and computing power requirements are lower than those of the DL methods, making it more suitable for deployment on resource-constrained EDs. Recent IoT ground sensing system projects have widely used ML anomaly detection algorithms, as summarised in Table VI.
@@
Table. VI presents an summary of the anomaly detection algorithms used in recent forest fire detection projects. In general, DL and ML approaches are widely used in these projects. Most of these initiatives employ a threshold limit, based on the z-score, regression, or LOF, for initial filtering. Upon triggering consecutive alerts, the ED transmits data to the server to perform fire detection or performs specific fire detection algorithms on the ED itself. Note that in this section, we have focused on a few anomaly detection techniques that are suitable for implementation on EDs or the gateway for wildfire detection. The selection of these methods was based on their computing power requirements, complexity, and whether they have been previously implemented.
Algorithm | Reference(s) | Pros | Cons |
Statistical Threshold | [96, 93, 100] | Easy to deploy | Inaccurate |
[186, 174, 172] | |||
Time Series Detection | [187] | Pattern-wise detection | Complexity in |
multivariate data | |||
ML Classfication | [97, 93, 100, 185] | Efficient multivariate | Require periodic |
[186, 191, 192] | data integration | re-training of models | |
DL Neural Networks | [81, 172, 174, 175] | Accurate | Resource requirement |
[193, 194, 187, 195, 191] | delay response |
@@ IV-A Sensors
As mentioned in Sec. II, gas/smoke sensors have different characteristics when it comes to wildfire detection. Some sensors have minimal cross-sensitivity, but their operating temperature and humidity ranges are limited. On tperformr hand, some sensors are highly accurate, but they take time to precisely measure the concentration of the target gases. Additionally, the cost and availability of these sensors, such as energy consumption, maintenance cost and size, are crucial considerations. In fact, numerous MoS sensors require different voltage levels to operate, such as the SGP30 and CCS811 which run at 1.8 volts. Using a linear voltage regulator on a 3.3 volt battery can lead to significant energy loss, and therefore a buck converter might be required to efficiently convert DC-to-DC voltage for these sensors. Although [196] indicates that the detection of fire flicker characteristics is not influenced by the types of combustible materials or the distance from the source, employing a typical IoT camera with a 62.2∘ horizontal and 48.8∘ vertical field of view to identify fires at operational distances up to 500 meters within a forest presents significant challenges. For example, to detect a fire from 500 meters away with a minimum required resolution of 16x16 pixels for reliable detection, the fire must be at least 5.03 meters wide and 6.72 meters tall, as per field of view calculations. In addition, it is essential to take into account not only the cost of the sensors, but also the operational costs associated with the equipment and needed for their operation and maintenance. A thorough assessment of these elements is necessary when selecting the right sensors for a particular purpose. Sensors can be delicate and need to be calibrated regularly, which can make them costly to use in a wildfire detection system. They may be exposed to extreme environmental conditions and may need to be replaced often. Besides, the pthat use edge computing.n of these sensors in the field also presents several challenges, including the establishment of a secure mounting site and the protection of the sensors against environmental hazards. These challenges pose obstacles in deploying sensors in remote regions, which are often the areas most vulnerable to wildfires.
@@ IV-B Communication link and Access Layer Control
Typically, IoT systems employ LPWAN protocols for communication between the end nodes and the gateway. These LPWAN communication protocols are characterised by ultra-narrowband or simple modulation/demodulation techniques, enabling low energy consumption during communication. The use of ultra-narrowband signals reduces the chance of packet collision, while wideband signals with distinctive patterns allow for easy differentiation from other signals, facilitating extended-range communication. Key players in the LPWAN domain include LoRaWAN, Sigfox, 6LoWPAN, NBIoT and LTE-M [198]. Although NBIoT and LTE-M operate within licensed spectrum bands, LoRaWAN together with Sigfox and 6LoWPAN utilise ISM-shared bands. A performance analysis comparing various wireless communication technologies, as shown in Fig. 6, demonstrates that LPWAN offers the longest range and the lowest energy consumption. However, it should be noted that LPWAN technologies exhibit the lowest data throughput. As a result, IoT system designers must prioritise the significance of data and determine which data to transmit and discard. Although most LPWAN technologies support packet fragmentation, this feature can exaggerate the likelihood of packet collisions and the overall packet loss rate within the system.
@@
In wireless communication, packet collision occurs when multiple packets are simultaneously transmitted to a receiver, making it impossible to differentiate between individual packets. In the absence of proper MAC layer control, IoT packets follow a random access communication model, with the packet successful rate denoted by , where is the traffic load measured by the total number of transmission attempts per frame time. If we assumed that all EDs operate at identical frame times within the system and transmit at uniform intervals, the limiting factor to achieve optimal packet success is defined by the expression , in accordance with the peak throughput characteristic of the ALOHA protocol, where represents the quantity of EDs, denotes the attempt rate per device, and signifies the frame time. To mitigate this issue, an MAC layer is responsible for regulating channel traffic, thereby ensuring successful packet reception. The MAC layer in most LPWAN technologies is typically predefined, and some employ random access or Listen Before Transmit (LBT) techniques [199, 200], while NB-IoT and LTE-M use the same MAC control as cellular networks due to their integration in the cellular communication system. [201] explores a modified LBT-type MAC control within LoRaWAN, demonstrating its ability to significantly reduce packet collisions and thus enhance support for numerous IoT EDs. Similarly, [202] presents simulation findings indicating that the average delay for LBT-LoRa is approximately 7 seconds compared to random access methods, further affirming the suitability of LBT for IoT communications. Another type of MAC layer control can be found in [203], where a custom Carrier-sense multiple access with collision avoidance (CSMA/CA) MAC layer was implemented in the LoRa communication system to improve the packet success rate. Nevertheless, the adoption of an in-house MAC layer control can lead to increased energy consumption for end devices due to requirements such as channel listening and synchronisation. An alternative approach is demonstrated in [204], where an in-house CSMA/CA MAC control for LoRa is introduced, resulting in an improvement in energy consumption. Overall, achieving a robust and energy-efficient communication link in IoT systems presents challenges that require thorough consideration of various factors, including gateway coverage influenced by channel loss and modulation schemes, channel interference resulting from traffic load, packet size and packet priorities, as well as quality of service and packet latency. In order to tackle these obstacles, the implementation of a custom MAC layer control emerges as a viable solution for IoT designers.
@@ IV-D Accuracy and Topology
As mentioned above, to minimise energy consumption in the ED, it is recommended to use energy-efficient detection algorithms. In cases where the decision-making process is based solely on the ED, simpler ML algorithms or statistical threshold detection methods can be employed. However, it is important to note that employing a straightforward detection algorithm, such as threshold detection or simplistic ML classification, can result in potential problems of increased sensitivity and a large number of false alarms. While specific measurement ranges are typically anticipated during unwanted fire incidents, other nonfire-related occurrences can also trigger sensor anomalies. Such anomalies can arise from animals or humans exhaling near the sensors. To improve the precision of fire detection systems, a viable solution is to integrate multiple detection aspects, such as combining vision-based detection and detection based on environmental monitoring sensors, to establish a remote sensing system. Nonetheless, this solution may incur additional energy consumption in each individual end device and additional delay in fire detection. As discussed in Subsection IV-B, the restricted number of EDs resulting from the limited bandwidth of the radio channel requires a throughout design of the system’s topology to achieve low latency and reduce packet collisions. In [206] and [207], innovative network topologies are introduced aimed at reducing energy use, enhancing system coverage, and preserving overall system efficiency. These models, while suitable for various IoT networks, assume that IoT gateways can be positioned anywhere in the field without affecting communication quality. In [208], a more practical approach is suggested, positioning gateways at the forest’s perimeter with sensors close to these gateways, showing an improvement in the energy efficiency of the system. However, the delay in detection with this topology model remains unexplored. Therefore, deciding on the optimal number of EDs, vision-based EDs, and the most suitable spacing and angular arrangement between EDs are essential for the IoT wildfire detection system. To address these topological issues, extensive measurement and analysis must be performed. In addition, reducing latency is crucial for prompt detection of wildfires, allowing faster responses when true fires occur. Similarly, it is important to minimise false alarms by evaluating the sensitivity of algorithms, particularly those using edge computing. Higher sensitivity improves a model’s ability to detect all true positive instances, which contributes to both reducing false negatives and accelerating response times to actual fires. In [209], a comparison of the sensitivity of various machine learning models suitable for edge computing is presented. The results show that Naive Bayes achieves the highest sensitivity at 96, followed by ANN at 95.5, DT at 93, and KNN at 93.9. In contrast, in [93], experiments reveal that the sensitivity of employing a statistical threshold is as low as 68.8.
@@
The challenges involved in implementing an IoT sensor system for early wildfire detection in remote areas can be illustrated by the trade-off between energy consumption, detection accuracy and delay. One of the main challenges is to conserve energy within the system. This can be achieved by using energy-friendly algorithms on the end devices and by sending decisions directly to the main server. However, this approach may not be able to guarantee high accuracy. Furthermore, ensuring detection accuracy requires a comprehensive sensing system that uses multiple aspects for detection and incorporates cross-validation techniques. However, this approach results in increased energy consumption and delay. Alternatively, in the pursuit of avoiding packet collision and achieving rapid response, only the detection results are transmitted to the server, which may lead to compromised accuracy. In [192], researchers performed an analysis of the performance of the IoT network employing both edge computing and cloud computing architectures. In the edge computing example, detection processes are executed on end devices, whereas in the cloud computing framework, these processes are centralised on the main server. The findings reveal that edge computing outperforms cloud computing with an average reduction of 70 ms in detection latency. Furthermore, it is observed that as the packet size increases, the detection latency in cloud computing proportionally escalates. The accuracy of machine learning models ranges between 98.7-99% for edge computing, and the presence of a large number of sensors introduces an additional processing delay of 2-5 seconds in the system. Although the detection accuracy for cloud computing is not specified, the study also demonstrates that a minimum interval between two consecutive readings is necessary to detect an anomaly. However, this interval may not be achievable with a single sensor reading. In other words, the ML-based anomaly detection model may require multiple successive sensor readings to identify an anomaly, and this problem can be exacerbated if spatial correlation is involved. Overall, there is a complex trade-off between energy consumption, detection accuracy and delay in implementing an IoT sensor system for wildfire detection in remote areas. There is no single solution that is optimal for all situations. The best approach will depend on several factors, including the size of the area to be monitored, the availability of resources and the risk of wildfires.
V Suggestions for Future Implementation